# Is It Really You? Exploring Biometric Verification Scenarios in Photorealistic Talking-Head Avatar Videos

### v1

@@ 6.3 Inter-Dataset Analysis
To assess generalization, we evaluate the performance of our biometric system across datasets. We train on one dataset (e.g., RAVDESS $\mathcal{D}_{\text{train}}$) and test on the other (e.g., CREMA-D $\mathcal{D}_{\text{test}}$), evaluating whether the system learns generalizable facial motion representations, or overfits to recording conditions, actor pools, or expression styles of a specific dataset. We follow the same protocol as described in Sec. 5. These inter-dataset experiments represent a more realistic real-world scenario.
@@ 7 Discussion
Our experiments show that facial motion-based biometric verification is feasible even in challenging avatar-mediated scenarios. Across multiple settings and datasets, our proposed biometric system has achieved AUC values surpassing 80% in the best cases. It is worth noting that our system exclusively focuses on landmarks alone, without using any facial appearance cues or conventional DeepFake detection features. This is a key design choice: in real avatar-mediated communication, all users share the same high-quality rendering pipeline, and appearance artifacts will not distinguish impostors from genuine users.
Our dataset generation strategy (using impostor avatars that perfectly match the victim’s appearance) forces the verification model to focus exclusively on behavioral facial motion patterns rather than static facial structure. If we had simply used the original real videos, the model could trivially rely on differences in facial geometry to verify identities. Such structural cues would be useless in a real attack scenario, where a stolen avatar would exactly replicate the victim’s face. By constructing impostor avatars, we ensure the model is trained on the realistic and challenging problem of detecting who is driving the avatar’s movements.
Compared to the only other published system for this task [16], which reports up to 87% AUC using a proprietary model trained on three datasets, our approach remains competitively despite using fewer data and GPU resources. Our biometric system is also lighter with reduced computational complexity, and provides explainability through the temporal attention weights revealing which frames and expressions the model considers most informative. This positive aspect is crucial, as described in the literature [4, 5].
One limitation of our biometric system is its exclusive reliance on landmark-based representations: while interpretable and efficient, they depend on the quality of the landmark estimator. We have used MediaPipe [11], a widely adopted and robust model, but inaccuracies can occur, especially under challenging poses or expressions, potentially reducing verification accuracy.
Another limitation is inherent to the avatar generation process itself. While GAGAvatar represents the state-of-the-art in one-shot talking-head synthesis [3], it does not always faithfully reproduce subtle or extreme facial expressions from the driver. Improving avatar generation fidelity, particularly in capturing fine-grained facial dynamics, could directly benefit biometric performance.
@@ 8 Conclusions
In this work, we have introduced the challenge of biometric verification in photorealistic talking-head avatar videos, where impostors can perfectly impersonate a victim’s appearance. We have proposed a lightweight, explainable system based solely on facial landmarks’ motion, using a spatio-temporal GCN with temporal attention-based pooling to learn discriminative behavioral signatures.
In addition to the proposed biometric system and code, we have also released a public standard benchmark of genuine and impostor avatar videos for this scenario. Our experiments have preliminary shown that facial motion patterns are effective biometric cues, demonstrating the potential of behavioral biometrics as a defense against avatar-based impersonation. Future work will be oriented to improve the performance through novel deep learning architectures [7, 9, 10] and loss functions [1, 8] and the evaluation of the biometric system under more challenging scenarios and avatar generation models [3, 12, 13, 21].

### v2

@@ 6.3 Inter-Dataset Analysis
We evaluate our biometric system across datasets to assess generalization. We train on one dataset (e.g., RAVDESS $\mathcal{D}_{\text{train}}$) and test on the other (e.g., CREMA-D $\mathcal{D}_{\text{test}}$), evaluating whether the system learns generalizable facial motion representations, or overfits to recording conditions, actor pools, or expression styles of a specific dataset. We follow the same protocol as in Sec. 5.
@@ 7 Discussion
Our experiments show that facial motion-based biometric verification is feasible even in challenging avatar-mediated scenarios. Across multiple settings and datasets, our proposed biometric system has achieved AUC values that surpass 80% in the best cases. It is worth noting that our system exclusively focuses on landmarks alone, without using any facial appearance cues or conventional DeepFake detection features. This is a key design choice: in real avatar-mediated communication, all users share the same high-quality rendering pipeline, and appearance artifacts will not distinguish impostors from genuine users.
Our dataset generation (using avatars of impostors that perfectly match the appearance of the victim) forces the verification model to focus exclusively on behavioral facial motion rather than static facial structure. If we had simply used the original real videos, the model could trivially rely on differences in facial geometry to verify identities. Such structural cues would be useless in a real attack scenario, where a stolen avatar would exactly replicate the victim’s face. By constructing impostor avatars, we ensure that the model is trained on the realistic and challenging problem of detecting who is driving the avatar’s movements.
Compared to the only other published system for this task [29], which reports up to 87% AUC using a proprietary model trained on three datasets, our approach remains competitive despite using fewer data and GPU resources. Our biometric system is also lighter with reduced computational complexity, and provides explainability through the temporal attention weights revealing which frames and expressions the model considers most informative. This positive aspect is crucial, as described in the literature [7, 8].
One limitation of our biometric system is its exclusive reliance on landmark-based representations: while interpretable and efficient, they depend on the quality of the landmark estimator. We have used MediaPipe [20], a widely adopted and robust model that has shown excellent tracking results in comparative studies [12], but inaccuracies can occur, especially under challenging poses or expressions, potentially reducing verification accuracy.
Another limitation is inherent in the avatar generation process itself. Although GAGAvatar represents the state-of-the-art in talking head one-shot synthesis [5], it does not always faithfully reproduce subtle or extreme facial expressions from the driver. Improving avatar generation fidelity, particularly in capturing fine-grained facial dynamics, could directly benefit biometric performance.
@@ 8 Conclusions
In this work, we have introduced the challenge of biometric verification in photorealistic talking head avatar videos, where impostors can perfectly impersonate the appearance of a victim. We have proposed a lightweight, explainable system based solely on facial landmarks’ motion, using a spatio-temporal GCN with temporal attention-based pooling to learn discriminative behavioral signatures.
In addition to the system code, we have also released a public standard benchmark of genuine and impostor avatar videos for this scenario. Our experiments have preliminary shown that facial motion patterns are effective biometric cues, demonstrating the potential of behavioral biometrics as a defense against avatar-based impersonation. Future work will be oriented to improve the performance through novel deep learning architectures [10, 16, 19] and loss functions [3, 26, 15] and the evaluation of the biometric system under more challenging scenarios and avatar generation models [5, 21, 22, 34]. Future work will also integrate other information beyond facial cues in multimodal approaches [28] with speech and other interaction signals [1, 2].